OpenStudy (anonymous):

Solve the following problem. Be sure to show all steps (V.E.S.T.) and work in order to receive full credit. The sum of three numbers is 26. The second number is twice the first and the third number is 6 more than the second. Find the numbers.?

OpenStudy (anonymous):

The sum of three numbers is 26." x+y+z = 26 "the second number is twice the first" y = 2x "the third number is 6 more than the second" z = 6+y You have the system of equations x + y + z = 26 2x - y = 0 -y + z = 6 Solve using Gauss-Jordan Elimination: x+y+z = 26 -3y-2z = -52 -y+z = 6 x+y+z = 26 y+(2/3)z = 52/3 -y+z = 6 x+(1/3)z = 26/3 y+(2/3)z = 52/3 (5/3)z = 70/3 x+(1/3)z = 26/3 y+(2/3)z = 52/3 z = 14 x = 4 y = 8 z = 14

OpenStudy (precal):

what is V.E.S.T.?

OpenStudy (precal):

just curious never seen that before

OpenStudy (anonymous):

Thank you @Ashley16

OpenStudy (anonymous):

Your welcome!